OpenStudy (anonymous):

Find the solution to –x2 – 2x – 5 = 0

OpenStudy (anonymous):

use the quadratic formula or solve by completing squares but first –x2 – 2x – 5 = 0 -(x2 +2x + 5 )=0 (x2 +2x + 5 )=0

OpenStudy (anonymous):

once again thanx

OpenStudy (anonymous):

using quadratic formula: [2 +- sq. root of 4 - 4(-1)(-5)] / -2 answer will be -1 - sq. root of 6 and -1 + sq. root of 6

OpenStudy (anonymous):

thanks bro